You are scheduled to receive annual payments of $8,600 for each of the next 25 years. The discount rate is 8.0 percent. What is the difference in the present value if you receive these payments at the beginning of each year rather than at the end of each year?

$7,344.25

$9,288.00

$8,312.76

$8,600.00

$7,763.93
Answer #1

At end of year:

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=8600[1-(1.08)^-25]/0.08

=8600*10.6747762

=91803.0753(Approx)

At beginning of year:

Present value of annuity due=Present value of annuity(1+rate)

=91803.0753*1.08

=99147.3213

Hence difference=99147.3213-91803.0753

=$7344.25(Approx).